I know that the equation to find the sum of the measures of the angles for a regular polygon is
(180)N-360 when N=Number of sides. But I do not understand why.
Why do you multiply N by 180 then subtract 360.
Thanks.

Found 2 solutions by Alan3354, Edwin McCravy:
Answer by Alan3354(69443)

I know that the equation to find the sum of the measures of the angles for a regular polygon is
(180)N-360 when N=Number of sides. But I do not understand why.
Why do you multiply N by 180 then subtract 360.
---------------
It's the same as 180*(n-2) = sum of angles
= 180*n - 360
-------------
If n = 3, 540 - 360 = 180 for a triangle.
If n = 4, 720 - 360 = 360 for a rectangle.
Don't argue with what works.
------------
The why is harder to explain.
============
I can do it on paper and scan it.
The other tutor didn't explain why there are 180 degs in a triangle.

Answer by Edwin McCravy(20054)

The other tutor said "Don't argue with what works".  I agree but

that's not what you did.  You aksed WHY it works, and you are to be

commended for asking that question, not criticized!!!

Here is the answer to your question:  

Any polygon can be divided up into triangles.  The number of triangles

is always two less triangles than the number of sides of the polygon.

For instance, take this 5-sided polygon (called a "pentagon", like the buiding 

in Washington DC): 



If you pick any vertex and draw the diagonals, 
like this:



the polygon is divided into three triangles (two less

than five).  The interior angles of the 3 triangles are

180° each.   So the sum of the interior angles of the

polygon is 3 times 180°.  That's

the number of sides, 5, minus 2, or 3 times 180°, and if the

number of sides is n, then that's

       (n - 2)180°

But as the other tutor explained, if you use the 
distributive principle, you get:

     n·180° - 2·180°

and that becomes

      180°n - 360°

Edwin